From a group of boys and girls, 14 girls leave first. Then the ratio of the number of girls to the number of boys becomes 1:2. After this, 43 boys leave the group. Now the boys and girls are equal in number. How many girls were there in the beginning?
Answer by richwmiller(17219)

(g-14)/b=1/2,
(b-43)/(g-14)=1
b = 86, g = 57